OpenStudy (ammarah):

Find the vertex and axis of symmetry of the graph of the function: f(x)=2x^2+8x

OpenStudy (mertsj):

Put the equation into this form: f(x)=2(x^2+4x) Then complete the square.

OpenStudy (ammarah):

how do u complete the square?

OpenStudy (mertsj):

Take half the coefficient of x and square it.

OpenStudy (ammarah):

y+4=2x^+4x+4

OpenStudy (mertsj):

y+8=2(x^2+4x+4)

OpenStudy (ammarah):

so i solve that to find the vertex?

OpenStudy (mertsj):

Put it into this form: y=(x-h)^2+k (h,k) is the vertex